OpenStudy (ja1):

Help me please. Fan and Medal! y = 3x2 + 12x + 5.

OpenStudy (ja1):

Identify the vertex for the graph of: \[y=3x^{2}+12x+5\]

OpenStudy (ja1):

@dumbcow @RadEn

OpenStudy (ja1):

@zepdrix

OpenStudy (ja1):

@tcarroll010 @whpalmer4

OpenStudy (raden):

the vertex of the that curve at point : (-b/2a, y(-b/2a))

OpenStudy (ja1):

Yes but how do you solve it since you do not have a GCF?

OpenStudy (raden):

y=3x^2+12x+5 known a=3, b=12, and c=5 x-vertex = -b/(2a) = -12/(2*3) = -12/6 = -2

OpenStudy (ja1):

OH RIGHT: \[\frac{ -b }{ 2a }\]

OpenStudy (raden):

then subtitute x=-2 into y=3x^2+12x+5to get y-vertec y = 3(-2)^2 + 12(-2) + 5 = 3(4) - 24 + 5 = -7 so, it is (-2,-7)

OpenStudy (ja1):

Right right, I forgot about that formula, thanks a bunch :D

OpenStudy (raden):

welcome :)